\begin{thebibliography}{99}
\bibitem{BFKL}  L. N. Lipatov, Sov. J. Nucl. Phys. ${\bf 23}$ (1976)
  642;\newline 
V. S. Fadin, E. A. Kuraev and L. N. Lipatov, Phys. Lett. ${\bf B60}$ (1975) 50;
Sov. Phys. JETP {\bf 44} (1976) 443; {\bf 45} (1977) 199;\newline
Ya. Ya. Balitsky and L. N. Lipatov, Sov. J. Nucl. Phys. ${\bf 28}$ (1978) 822.

\bibitem{NLO}  V. S. Fadin, L.N. Lipatov, Phys. Lett. ${\bf {B429}}$ (1998)
127;\newline
G. Camici and M. Ciafaloni, Phys. Lett. {\bf B430} (1998) 349.;\newline
A. V. Kotikov, L. N. Lipatov, Nucl. Phys. {\bf B582} (2000); hep-ph/0208220.

\bibitem{DGLAP}  V. N. Gribov and L. N. Lipatov, Sov. J. Nucl. Phys. {\bf 18}
(1972) 438, 675;\newline
L. N. Lipatov, Sov. J. Nucl. Phys. {\bf 20} (1975) 93;\newline
G. Altarelli and G. Parisi, Nucl. Phys. {\bf B126} (1977) 298; \newline
Yu. L. Dokshitzer, Sov. Phys. JETP {\bf 46} (1977) 641.

\bibitem{report}  L. N. Lipatov, {\it Pomeron in QCD}, in `Perturbative
QCD', ed. by A. N. Mueller, World Scientific, 1989; {\it `Small-x physics in
the perturbative QCD'}, Phys. Rep., {\bf 286} (1997) 132.

\bibitem{conf}  L. N. Lipatov, Sov. Phys. JETP {\bf 63} (1986) 904.

\bibitem{holom}  L. N. Lipatov, Phys. Lett. ${\bf {B251}}$ (1990) 284.

\bibitem{photons}  S. J. Brodsky, V. S. Fadin, V. T. Kim,
  L. N. Lipatov, G. V. Pivovarov, JETP Letters {\bf B70} (1999) 155;
  {\bf B76} (2002) 306.

\bibitem{T} See for example, J. I. Kapusta, {\it Finite Temperature Field
Theory}, Cambridge
Univ. Press, 1989; M. Le Bellac, {\it Thermal Field Theory},
Cambridge University Press, 1996.

\bibitem{QGP}  {\it Quark-Gluon Plasma 3}, edited by R. C. Hwa and X. N. Wang,
World Scientific, September 2003.
E. V. Shuryak, `The QCD vacuum, hadrons and superdense matter',
World Sci. Lect. Notes Phys. 8: 1-401, 1988.
H. Satz, Rept. Prog. Phys. 63, 1511 (2000). See also, E. V. Shuryak,
I. Zahed, hep-ph/0307267.

\bibitem{BKP}  J. Bartels, Nucl. Phys. ${\bf {B175}}$ (1980) 365;\newline
J. Kwiecinski and M. Prascalowicz, Phys. Lett. ${\bf {B94}}$ (1980) 413.

\bibitem{intodd}  L. N. Lipatov, Phys. Lett. ${\bf {B309}}$ (1993) 394.

\bibitem{brown}  L. N. Lipatov, Proceedings of the Fifth Blois Workshop,
edited by H.M. Fried, K.Kang, C-I Tan, June 1993, World Scientific.

\bibitem{integr}  L. N. Lipatov, hep-th/9311037, Padova preprint
DFPD/93/TH/70, unpublished.

\bibitem{Heis}  L. N. Lipatov, Sov. Phys. JETP Lett. ${\bf {59}}$ (1994) 571.

\bibitem{FadKor}  L. D. Faddeev, G. P. Korchemsky, Phys. Lett. {\bf {342}}
(1995) 311.

\bibitem{JW}  R. Janik and J. Wosiek, Phys. Rev. Lett. ${\bf 79}$ (1997)
2935; {\bf 82 }(1999) 1092.

\bibitem{dual}  L. N. Lipatov, Nucl. Phys. {\bf B548} (1999) 328.

\bibitem{new}  J. Bartels, L. N. Lipatov, G. P. Vacca, Phys.Lett. {\bf
B477} , 178 (2000).

\bibitem{hect}  H. J. de Vega and L. N. Lipatov, Phys. Rev. {\bf D64},
114019 (2001); {\bf D66}, 074013 (2002).

\bibitem{dkm}  A. Derkachev, G. Korchemsky, A. Manashov, Nucl. Phys. {\bf 
B617 } (2001) 375.

\bibitem{BalKov}  I. Balitsky, Phys. Rev. {\bf D60,\ }014020 (1999);\newline
Y. V. Kovchegov, Phys. Rev. {\bf D61,\ }074018 (2000).

\bibitem{eff}  L. N. Lipatov, Nucl. Phys. {\bf B452,} (1995) 369.

\bibitem{bh} See for example,
N. D. Birrell, P. C. W. Davis, `Quantum Fields in Curved Space',
Cambridge U. Press, 1982 and references therein.
S. W. Hawking, Comm. Math. Phys. 43, 199 (1975).
N. Sanchez, Phys. Lett. B87, 212 (1979), Phys. Rev. D24, 2100 (1981).

\end{thebibliography}

\documentstyle[12pt]{article}
\begin{document}
\begin{titlepage}

\centerline{\bf BFKL Pomeron at non-zero temperature} 

\centerline{\bf and integrability of the Reggeon dynamics in multi-colour QCD}

\date{\today}

\begin{center}
H. de Vega \\
Laboratoire de Physique Th\'eorique et Hautes Energies,\\
Universt\'e Pierre et Marie Curie, Paris VI et Universt\'e Denis
Diderot, Paris VII, BP 126,\\
4, place Jussieu, F-75252 Paris cedex 05, France.\\
Laboratoire Associ\'e au CNRS UMR 7589.\\
\end{center}
\begin{center}
L.N. Lipatov {$^{*}$\\
Theoretical Physics Department,\\
Petersburg Nuclear Physics Institute,\\
Gatchina, 188300, St. Petersburg, Russia}
\end{center}
\vskip 15.0pt

\centerline{\bf Abstract}

\noindent
We consider the QCD scattering amplitudes at high energies $\sqrt{s}$
and fixed momentum transfers $\sqrt{-t}$ in the leading logarithmic
approximation at a non-zero temperature $T$ in the $t$-channel.  It is
shown that the BFKL Hamiltonian has the property of holomorphic
separability. The Pomeron wave function for arbitrary $T$ is calculated
using an integral of motion. In multi-colour QCD, the
holomorphic Hamiltonian for $n$-reggeized gluons at temperature $T$ is
shown to coincide with the local Hamiltonian of an integrable
Heisenberg model and can be obtained from the $T=0$ Hamiltonian by
an unitary transformation.
We discuss the wave functions and the spectrum of intercepts for the
colourless reggeon states.  \vskip 3cm

\hrule

\vskip 3cm

\noindent

\noindent $^{*}$ {\it Supported by the grant
INTAS 2000-366} \vfill

\end{titlepage}

1. In QCD the scattering amplitudes $A(s,t)$ in Regge kinematics for high
energies $2E=\sqrt{s}$ and fixed momentum transfers $q=\sqrt{-t}$ are
obtained in the leading logarithmic approximation $\alpha _{s}\ln s\sim
1,\,\alpha _{s}=\frac{g^{2}}{4\pi }\rightarrow 0$ ($g$ is the QCD coupling
constant) by summing the largest contributions $\sim \left( \alpha _{s}\ln
s \right) ^{n}$ to all orders of perturbation theory within
the approach of Balitsky, Fadin, Kuraev and Lipatov (BFKL) \cite{BFKL}. The
BFKL Pomeron in the $t$-channel turns out to be a composite state of two
reggeized gluons (it is valid also in the next-to-leading approximation
\cite{NLO}). Its wave function $\Psi (\vec{\rho}_{1},\vec{\rho}_{2})$ satisfies
the Schr\"{o}dinger equation in the two dimensional impact-parameter space
$\vec{\rho}$
\begin{equation}
E\,\Psi (\vec{\rho}_{1},\vec{\rho}_{2})=H_{12}\,
\Psi (\vec{\rho}_{1},\vec{\rho}_{2})\;.
\end{equation}
The intercept $\Delta $ of the Pomeron, related to the high energy
asymptotics $\sigma _{t}\sim s^{\Delta }$ of the total cross-section, is
proportional to the ground state energy $E$
\[
\Delta =-\frac{\alpha _{s}N_{c}}{2\pi }\,E\;.
\]
The kinetic part $H_{kin}=\ln \,|p_{1}|^{2}+\ln \,|p_{2}|^{2}$ of $H_{12}$
is a sum of two gluon Regge trajectories and its potential part $H_{pot}$ is
related by a similarity transformation to the two-dimensional Green
function $\ln \,|\rho _{12}|^{2}$, where $\rho _{12}=\rho _{1}-\rho _{2}$.
[We introduced here the complex coordinates $\rho _{r}=x_{r}+iy_{r}$ and the
corresponding momenta $p_{r}=i\partial _{r}$].

The BFKL equation is used for the description of the deep-inelastic
lepton-hadron scattering together with the DGLAP equation \cite{DGLAP} (see
for example \cite{report}). It is invariant under the M\"{o}bius
transformations \cite{conf}
\[
\rho _{r}\rightarrow \frac{a\rho _{r}+b\rho _{r}}{c\rho _{r}+d\rho _{r}}
\]
with arbitrary complex parameters $a,b,c,d$ and $H_{12}$ has the property of
holomorphic separability (see \cite{report} \cite{holom})
\begin{equation}\label{HT0}
H_{12}=h_{12}+h_{12}^{\ast }\,,\,\,h_{12}=\sum_{r=1}^{2}\left[ \ln \,p_{r}+
\frac{1}{p_{r}}\,\ln \,(\rho _{12})\,\,p_{r}-\psi (1)\right] \;  ,
\label{hamil}
\end{equation}
where $\psi (z)=\Gamma ^{\prime }(z)/\Gamma (z)$.

The wave functions $\Psi $ belong to the principal series of unitary
representations of the M\"{o}bius group with conformal weights $
m=1/2+i\nu +n/2,\,\,\widetilde{m}=1/2+i\nu -n/2$ expressed in terms of the
anomalous dimension $\gamma =1+2i\nu $ and the integer conformal spin $n$
for the local gauge-invariant operators \cite{conf}. The conformal weights
are related to the eigenvalues $m(m-1)$, $\widetilde{m}(\widetilde{m}-1)$ of
the Casimir operators $M^{2}$ and $M^{2\ast }$, where
\[
M^{2}=\left( \sum_{r=1}^{2}M_{3}^{(r)}\right) ^{2}+\frac{1}{2}\left(
\sum_{r=1}^{2}M_{+}^{(r)}\,\sum_{s=1}^{2}M_{-}^{(s)}+
\sum_{r=1}^{2}M_{-}^{(r)}\,\sum_{s=1}^{2}M_{+}^{(s)}\right) =\rho
_{12}^{2}\; p_{1}\; p_{2}\;.
\]
Here $\vec{M}^{(r)}$ are the M\"{obius} group generators
\[
M_{3}^{(r)}=\rho _{r} \, \partial _{r}\,,\,\,\,M_{+}^{(r)}=\partial
_{r}\,,\,\,\,M_{-}^{(r)}=-\rho _{r}^{2} \; \partial _{r}
\]
and $\partial _{r}=\partial /\partial \rho _{r}$.

The eigenfunctions of $H_{12}$ can be considered as the three-point
functions of a two-dimensional conformal field theory and have the property of
holomorphic factorization \cite{conf},
\begin{equation}\label{fmm}
f_{m,\widetilde{m}}(\overrightarrow{\rho _{1}},\overrightarrow{\rho _{2}};
\overrightarrow{\rho _{0}})=\left\langle 0\right| \,\varphi
(\overrightarrow{
\rho _{1}})\,\varphi (\overrightarrow{\rho _{1}})\,O_{m\widetilde{,m}}(
\overrightarrow{\rho _{0}})\,\left| 0\right\rangle =\left( \frac{\rho
_{12}}{\rho _{10}\,\rho _{20}}\right) ^{m}\left( \frac{\rho _{12}^{\ast }}{\rho
_{10}^{\ast }\,\rho _{20}^{\ast }}\right) ^{\widetilde{m}}\,.
\end{equation}
One can calculate the energy putting this Ansatz in the BFKL
equation\cite{BFKL}
\begin{equation}
E_{m,\widetilde{m}}=\varepsilon _{m}+\varepsilon _{\widetilde{m}}
\quad , \quad \varepsilon _{m}=\psi (m)+\psi (1-m)-2\psi (1)\; .
\end{equation}
The minimum of $E_{m,\widetilde{m}}$ is obtained at $m=\widetilde{m}=1/2$
leading to a large intercept $\Delta =4 \; \frac{\alpha _{s}}{\pi
}N_{c} \, \ln 2$ of the BFKL Pomeron. In the next-to-leading
approximation the intercept is comparatively small ($\Delta \sim 0.2$
for the QCD case) \cite{photons}.

2. On the other hand, now a significant interest is devoted to the
quark-gluon plasma (QGP) generation in heavy nucleus
collisions (see for example \cite{QGP}). Current theoretical
understanding suggests that the QGP thermalizes via parton-parton
scattering. The QGP is understood to  cool down by hydrodynamic
expansion till the temperature reaches the hadronization scale $\sim 160$MeV.
One interesting phenomena is the suppression of the $\psi $-meson
production in the heavy nucleus collisions due to the
disappearance of the confining potential between $q$ and $\bar{q}$ at
high temperature \cite{QGP}. 
A similar effect should exist for glueballs
constructed from gluons. Because the Pomeron is considered as a composite
state of reggeized gluons, the influence of the temperature on its
properties is of great interest. In this paper we construct the BFKL
equation at temperature $T$ in the center-mass system of the $t$
-channel (where $\sqrt{t}=2\epsilon $) and investigate the integrability
properties of the BFKL dynamics in a thermostat for composite states of $n$
reggeized gluons in multi-colour QCD.

Let us consider the Regge kinematics in which the total particle
energy $\sqrt{s}$ is asymptotically large in comparison with the
temperature $T$. In this case one can neglect
the temperature effects in the propagators of the initial and intermediate
particles in the direct channels $s$ and $u$. But the momentum transfer
$|q|$ is considered to be of the order of $T$ (note, that $q_{\mu }$
is the vector orthogonal to the initial momenta $q_{\mu }\approx
q_{\mu }^{\perp }$ ). As it is well known \cite{T}, the particle wave
functions $\psi (x_{\mu })$ at temperature $T$ are periodic in the
euclidean time $\tau =i\,t$ with period $1/T$.

We introduce the temperature $T$ in the center of mass frame of the
$t$-channel. Thus, the euclidean energies of the intermediate
gluons in the $t$-channel become quantized as
$$
k_{4}^{(l)}=2\pi l\,T \; .
$$
In the $s$-channel the
invariant $t$ is negative and therefore the analytically continued 4-momenta
of the $t$-channel particles can be considered as euclidean vectors. It
means, that at temperature $T$, the wave functions
for virtual gluons are periodic functions of the 
holomorphic impact-parameter
$\rho =x+iy$ 
with imaginary period $\frac{i}{T}$. Also,
the canonically conjugated momenta $p$ have their imaginary part quantized,
\begin{equation}
\rho \rightarrow \rho +\frac{i}{T} \quad , \quad p=\mbox{Re}\, p+ 
\pi\, i \, l \, T \; .
\end{equation}
with integer $l$ (note that $p=(p_1+ip_2)/2 $).

It is convenient to rescale the transverse coordinates and corresponding
momenta as follows
\[
\rho \rightarrow \frac{1}{2\pi \,T} \; \rho \quad , \quad
p\rightarrow 2\pi \,T \, p \;.
\]
In these dimensionless variables one obtains
\[
0<\mbox{Im} \, \rho <2\pi \quad , \quad \mbox{Im} \, p= \frac{l}{2} \,.
\]
The calculation of the Regge trajectory $1+\omega (t)$ of the gluon 
at temperature $T$ in the t-channel,
in one-loop approximation reduces to the integration over the
real part $k_{1}$ of the transverse momentum $\vec{k}_{\perp}$
of the virtual gluon and to the summation over its imaginary part
$k_{2}=l$. In such a way we obtain
the following result for the trajectory having the separability
property [cf. \cite{holom}],
\[
\omega (-\vec{q}^{2})=-\frac{g^{2}}{8\pi ^{2}} \; N_{c} \; \Omega (-\vec{q}
^{2})\quad , \quad  \Omega (-\vec{q}^{2})=\Omega (q)+\Omega (q^{\ast }) \; .
\]
Here,
\[
\Omega (q)=\frac{\pi T}{\lambda }+\frac{1}{2}\,\left[ \psi (1+iq)+\psi
(1-iq)-2\psi (1)\right] \; ,
\]
where we regularized the infrared divergence for the zero mode $l=0$
introducing a mass $\lambda $ for the gluon (see \cite{BFKL}).

A similar  divergence appears in the Fourier transformation $G(
\overrightarrow{\rho }_{12})$ of the effective gluon propagator $(\vec{k}
_{\perp }^{2}+\lambda ^{2})^{-1}$ contained in the product of the effective
vertices $q_{1}k^{-1}q_{2}^{\ast }$ for the production of a gluon
with momentum $k_{\mu }$ (cf. \cite{report})
\[
G(\overrightarrow{\rho }_{12})=-\frac{\pi \,T}{\lambda }+\ln \left( 2\sinh
\frac{\rho _{12}}{2}\right) +\ln \left( 2\sinh \frac{\rho _{12}^{\ast }}{2}
\right) \; .
\]
Therefore, the divergence at $\lambda \rightarrow 0$ cancels in the
sum of kinetic and potential contributions to the BFKL equation and the
Hamiltonian $H_{12}$ for the Pomeron in a thermostat has the property of
holomorphic separability with the holomorphic Hamiltonian given below [cf.
eq.(\ref{hamil})]
\begin{equation}
h_{12}=\sum_{r=1}^{2}\left[ \frac12 \; \psi (1+ip_{r})+
\frac12 \; \psi(1-ip_{r}) +
\frac{1}{p_{r}}\,\ln \left( 2\sinh \frac{\rho_{12}}{2}\right) \,p_{r}-
\psi(1)\right] \; .  \label{hamilT}
\end{equation}

3. The Hamiltonian (\ref{hamilT}) is a periodic function of $\rho _{12}$.
Therefore its eigenfunctions are quasi-periodic functions of this
variable. The behaviour  of $h_{12}$ for small $\rho _{12}$
corresponds to
the low temperature regime. Hence, the eigenfunctions of $h_{12}$
behave for $  \rho_{12} \to 0 $ as the holomorphic part of the
zero-temperature wave functions eq.(\ref{fmm}),
\begin{equation}\label{ro0}
\Psi_m (\rho _{12}) \buildrel{ \rho_{12} \to 0}\over= \rho_{12}^{m}
\end{equation}
Notice that $ \Psi_{1-m} (\rho _{12})$ is an eigenfunction too.

We find the small-$T$ expansion of $h_{12}$ near its singularities $\rho
_{12}=2\pi il$. For example, for small $\rho _{12}$ and large
$p_{1},\,p_{2}$ we have
\[
h_{12} = h_{12}^{0}+\sum_{k=1}^{\infty} \frac{B_{2k}}{2k} \sum_{r=1,2}\left[
\frac{(-1)^{k+1}}{p_{r}^{2k}}+\frac{1}{p_{r}}\,
\frac{\rho_{12}^{2k}}{(2k)!} \,\,p_{r}\right] \; ,
\]
where $h_{12}^{0}$ is the holomorphic BFKL Hamiltonian at a zero
temperature [given by eq.(\ref{HT0})] and the $B_{2k}$ are the
Bernoulli numbers. This
representation for $h_{12}$ permits us to find the small-$T$ expansion of its
eigenfunctions. For example, at a vanishing momentum transfer $Q=p_{1}+p_{2}$
we obtain for the eigenfunction $\Psi_m (\rho _{12})$,
\[
\Psi_m (\rho _{12})=\rho_{12}^{m}\left[
  1-\frac{1}{24}\frac{m(m-1)}{2m+1} \;
  \rho_{12}^{2}+\frac{1}{5760}\frac{m(m-1)(5m^{2}+7m+6)}{(2m+1)(2m+3)}
  \; \rho_{12}^{4}+ {\cal O}(\rho_{12}^6) \right] .
\]
These eigenfunctions are parametrized by the conformal weight $m$.

On the other hand, from the above expansion it is possible to verify, that
the holomorphic Hamiltonian has the non-trivial integral of motion:
\begin{equation}\label{A}
A=4\sinh ^{2}\frac{\rho _{12}}{2}\,\,p_{1}\,p_{2}\quad , \quad
[A,h_{12}]=0\,.
\end{equation}
Therefore, instead of solving the Schr\"{o}dinger equation we can
search for the  eigenfunctions of the operator $A$. For non-zero $Q$
one can write the holomorphic wave function as a product of a plane wave
depending on $R=(\rho _1+\rho_2)/2$ times a
solution of the following equation for the relative motion of two gluons
\[
\left[ \frac{Q^{2}}{4}+\frac{\partial ^{2}}{\partial \rho^2}\right]
\; \Psi (\rho ,Q)=\frac{m(m-1)}{4 \sinh ^{2}\frac{\rho}{2}}\; \;\Psi (\rho ,Q)
\quad , \quad
\rho =\rho_{12}\,,\,\,t=-4\left| Q\right| ^{2} \; .
\]
The two independent solutions of the above differential equation can be
expressed in terms of hypergeometric functions
\begin{equation}
\Psi_{1}^{(m)}(\rho ,\,Q)=e^{\frac{i}{2} Q \, \rho} \; \left(e^{\rho}
- 1 \right)^m \; F\left(iQ+m,m;2m;1 - e^{\rho}\right)
\quad , \quad \Psi_{2}^{(m)}(\rho ,\,Q) \equiv \Psi _{1}^{(1-m)}(\rho
,\,Q) \; .
\end{equation}
For $\rho \rightarrow 0$ eq.(\ref{ro0}) holds and the singularities of
$\Psi^{(r)}(\rho ,\,Q)$ at $1 - e^{\rho}=1$ and $ 1 - e^{\rho}=\infty
$ correspond to the points $\rho =-\infty $ and $\rho =\infty $, respectively.

The analytic continuation of $\Psi ^{(r)}$ along the imaginary axes from $
\rho =0$ to $\rho =2\pi i$ is equivalent to the continuation of these
eigenfunctions in a circle passed in a clock-wise direction around the
singularity at $\rho =-\infty $. The monodromy matrix expressing the
analytically continued  solutions in terms of the initial ones can be
easily calculated.

The Pomeron wave functions can be written as a
bilinear combination of  holomorphic and anti-holomorphic eigenfunctions $\Psi
^{(r)}(\rho ,\,Q)$ and $\Psi^{(r)}(\rho^{\ast } ,\,Q^{\ast })$. The
property of single-valuedness in the cylinder topology
corresponding to the periodicity on the boundaries of the strip
$0<\mbox{Im} \,  \rho_{12}<2\pi $ is easily imposed to such Pomeron
wave function using  the monodromy matrix for $\Psi ^{(r)}(\rho
,\,Q)$. The resulting wave function can be written as,
\begin{equation}\label{hip}
\Psi ^{(m,\widetilde{m})}(\vec{\rho},\,\vec{Q})=\chi _{1}^{(m)}(\rho
,\,Q)\; \chi _{1}^{(\widetilde{m})}
(\rho^{\ast },\,Q^{\ast }) - (-1)^N
\; \chi _{2}^{(m)}(\rho ,\,Q)\;\chi _{2}^{(\widetilde{m})}(\rho ^{\ast
},\,Q^{\ast })\,,
\end{equation}
where,
$$
\chi _{1}^{(m)}(\rho ,\,Q) = 2^{1-2\,m} \;
\frac{\Gamma\left(m+iQ\right)}{\Gamma\left(m+\frac12\right)} \;
\Psi_{1}^{(m)}(\rho ,\,Q) \; , \; \chi _{2}^{(m)}(\rho ,\,Q) =
\chi_{1}^{(1-m)}(\rho ,\,Q) 
$$
and $ N = 2 \, \mbox{Im} \, Q $ is an integer.

4. The Pomeron wave function can be constructed directly in coordinate
space. For this purpose we use the conformal transformation
\begin{equation}\label{exp}
\rho _{r}=\ln \,\rho _{r}^{\prime }
\end{equation}
and the integral of motion eq.(\ref{A}) becomes
\[
A=-(\rho _{12}^{\prime })^{2}\frac{\partial }{\partial \rho
_{1}^{\prime }}\, \frac{\partial }{\partial \rho _{2}^{\prime }}\,.
\]
Thus, $A$ coincides in the variables $\rho _{r}^{\prime }$
with the Casimir operator of the conformal group whose eigenfunctions are
well known (see \cite{conf}). Thus, the Pomeron wave function at non-zero
temperature having the property of single-valuedness and periodicity
takes the form
\begin{equation}
\Psi ^{(m,\widetilde{m})}(\vec{\rho _{1}},\,\vec{\rho _{2}},\,\vec{\rho
_{0}})=\left( \frac{\sinh \frac{\rho _{12}}{2}}{2\sinh \frac{\rho _{10}}{2}
\,\sinh \frac{\rho _{20}}{2}}\right) ^{m}\,\left( \frac{\sinh \frac{\rho
_{12}^{\ast }}{2}}{2\sinh \frac{\rho _{10}^{\ast }}{2}\,\sinh \frac{\rho
_{20}^{\ast }}{2}}\right) ^{\widetilde{m}}\,.
\end{equation}
The orthogonality and completeness relations for these functions can be
easily obtained from the analogous results for $T=0$ (see \cite{conf})
using the above conformal transformation. These wave functions are 
proportional to the Fourier transformation
of the wave functions $\Psi ^{m,\widetilde{m}}(\vec{\rho},\vec{Q})$.

Moreover, the pair BFKL Hamiltonian $h_{12}$ can be expressed in terms of
the BFKL Hamiltonian at zero temperature in the new variables
\begin{equation}
h_{12}=\ln (p_{1}^{\prime }\,p_{2}^{\prime })+\frac{1}{p_{1}^{\prime }}
\,\log (\rho _{12}^{\prime })\,\,p_{1}^{\prime }+\frac{1}{p_{2}^{\prime }}
\,\log (\rho _{12}^{\prime })\,\,p_{2}^{\prime }-2\psi (1)\,,
\end{equation}
where $p_{r}^{\prime }=i\frac{\partial }{\partial \rho _{r}^{\prime
}}$. In the course of the derivation the 
following operator identity (see \cite{report})
\[
\frac{1}{2}\left[ \psi \left(1+z\frac{\partial }{\partial z}\right)+\psi
  \left(-z\frac{ \partial }{\partial z}\right)\right] =\ln \,z+\ln
  \,\frac{\partial }{\partial z}
\]
was used to transform the kinetic part as well as  properties of the
$\psi $-function.

In summary, the exponential mapping eq.(\ref{exp}) which in dimensional
variables takes the form
$$
\rho' = \frac{1}{2\pi \, T} \; e^{2\pi \, T \; \rho} \; ,
$$
maps the reggeon dynamics from zero temperature to temperature
$T$. This mapping explicitly exhibits a periodicity 
$\rho \rightarrow \rho +\frac{i}{T}$ for a thermal state. 
It must be noticed that such class
of mappings are known to describe thermal situations for
quantum fields in accelerated frames and in black hole backgrounds\cite{bh}.

\bigskip

5. As it is well known \cite{BKP}, the BFKL equation at $T=0$ can be
generalized to composite states of $n$ reggeized gluons. In
the multi-colour limit $N_{c}\rightarrow \infty $ the BKP equations are
significantly simplified thanks to their conformal invariance
\cite{conf},  holomorphic separability \cite{holom} and integrals of
motion \cite{intodd}. The generating function for the holomorphic
integrals of motion
coincides with the transfer matrix for an integrable lattice spin model
\cite{brown} \cite{integr}. The transfer matrix is the trace of the
monodromy matrix
\[
t(u)=L_{1}(u)\,L_{2}(u)...L_{n}(u)\,,
\]
satisfying the Yang-Baxter equations \cite{integr}. The integrability of the
$n$-reggeon dynamics in multi-colour QCD is valid also at non-zero
temperature $T$, where, according to the above arguments we should
take the $L$-operator in the form
\[
L_{k}=\left(
\begin{array}{cc}
u+p_{k} & e^{-\rho _{k}}\,p_{k} \\
-e^{\rho _{k}}\,p_{k} & u-p_{k}
\end{array}
\right) \,.
\]
In particular, the holomorphic Hamiltonian is the local Hamiltonian of the
integrable Heisenberg model with the spins being unitarily transformed
generators of the M\"{o}bius group (cf. \cite{Heis} \cite{FadKor})
\[
M_{k}=\partial _{k} \quad ,  \quad M_{+}=e^{-\rho _{k}}\,\partial
_{k} \quad ,  \quad M_{-}=-e^{\rho _{k}}\,\partial _{k}\; .
\]
Because the Hamiltonian at  non-zero temperature can be obtained by an
unitary transformation from the zero temperature Hamiltonian, the
spectrum of the intercepts
for multi-gluon states is the same as for zero temperature \cite{JW}-
\cite{dkm} and the wave functions of the composite states can be calculated
by the substitution $\rho _{k}\rightarrow e^{\rho _{k}}$. 

Furthermore, the non-linear Balitsky-Kovchegov equation \cite{BalKov}
can be generalized to the case of non-zero temperature as follows,
\begin{equation}
\frac{\partial N_{\vec{\rho}_{1},\vec{\rho}_{2}}}{\partial
  Y}=\bar{\alpha}
_{s}\int \frac{d^{2}\rho _{0}}{2\pi }\,\frac{\left| \sinh \frac{\rho
_{12}}{2}\right| ^{2}}{4\left| \sinh \frac{\rho _{10}}{2}\right|
  ^{2}\left| \sinh 
\frac{\rho _{20}}{2}\right| ^{2}}\,\left( N_{\vec{\rho}_{1},\vec{\rho}
_{0}}+N_{\vec{\rho}_{2},\vec{\rho}_{0}}-N_{\vec{\rho}_{1},\vec{\rho}_{2}}-N_
{\vec{\rho}_{1},\vec{\rho}_{0}}\,N_{\vec{\rho}_{2},\vec{\rho}_{0}}\right) \,,
\end{equation}
where $N_{\vec{\rho}_{1},\vec{\rho}_{2}}$ is the amplitude of finding \ a
dipole with the impact parameters $\vec{\rho}_{1}$ and $\vec{\rho}_{2}$ in a
hadron and the integration over $\rho _{0}$ is performed over the strip $0<
 \mbox{Im}\,\rho _{0}<2\pi $. Note, however, that in this equation  one
takes into account only
fan diagrams for the Pomeron interactions  
among all possible diagrams for reggeized gluons 
appearing in  the high energy effective action \cite{eff}.

We thank G.\thinspace S. Danilov and V. N. Velizhanin for helpful
discussions.

\end{document}